Express the function h(x)=\frac{1}{x + 9} in the form f \circ g. If g(x)=x + 9, find the function f(x).
Answer by jim_thompson5910(35256)

If f(x) = 1/x and g(x) = x+9, then (f o g)(x) = 1/(x+9)


You basically start with 1/x and replace x with x+9 to find f o g